OpenStudy (anonymous):

Choose the point on the terminal side of 135°. ANSWERS: (-1, sqrt3) (-sqrt3, 1) (-1, 1) (sqrt3, -1)

OpenStudy (anonymous):

@satellite73 , @amistre64

OpenStudy (anonymous):

look at the unit circle on the last page of the attached cheat sheet

OpenStudy (anonymous):

find the point on the unit circle corresponding to the angle of 135 degrees

OpenStudy (anonymous):

what do you see for the coordinates?

OpenStudy (anonymous):

i have one its-sqrt2/2, sqrt2/2

OpenStudy (anonymous):

i realize you do not see \((-1,1)\) but instead you see \((-\frac{\sqrt{2}}{2},\frac{\sqrt{2}}{2})\)

OpenStudy (anonymous):

so it (-1, 1)

OpenStudy (anonymous):

they are the same except one is negative and the other is positive that means if you go further out, you will get to the point \((-1,1)\)

OpenStudy (anonymous):

yes you can also use the "reference" triangle 

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

if you put in in the right place on the coordinate axis, you will get an angle of 135 degrees 

OpenStudy (anonymous):

yw